# Another look at inference after prediction

### v2

@@ C.1 Proof of Theorem 2
we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$. Thus,
${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$

### v3

@@ C.1 Proof of Theorem 2
we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$. Thus,
${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$